# OpenDlign: Enhancing Open-World 3D Learning with Depth-Aligned Images

### v2

@@ Abstract
Recent open-world 3D representation learning methods using Vision-Language Models (VLMs) to align 3D data with image-text information have shown superior 3D zero-shot performance. However, CAD-rendered images for this alignment often lack realism and texture variation, compromising alignment robustness. Moreover, the volume discrepancy between 3D and 2D pretraining datasets highlights the need for effective strategies to transfer the representational abilities of VLMs to 3D learning. In this paper, we present OpenDlign, a novel open-world 3D model using depth-aligned images generated from a diffusion model for robust multimodal alignment. These images exhibit greater texture diversity than CAD renderings due to the stochastic nature of the diffusion model. By refining the depth map projection pipeline and designing depth-specific prompts, OpenDlign leverages rich knowledge in pre-trained VLM for 3D representation learning with streamlined fine-tuning. Our experiments show that OpenDlign achieves high zero-shot and few-shot performance on diverse 3D tasks, despite only fine-tuning 6 million parameters on a limited ShapeNet dataset. In zero-shot classification, OpenDlign surpasses previous models by 8.0% on ModelNet40 and 16.4% on OmniObject3D. Additionally, using depth-aligned images for multimodal alignment consistently enhances the performance of other state-of-the-art models.
@@ 1 Introduction
3D understanding is increasingly pivotal in real-world applications like augmented/virtual reality [1, 2], autonomous vehicles [3, 4], and robotics [5, 6]. Traditional 3D models [7, 8, 9, 10, 11] are closed-world, recognizing only pre-defined categories and struggling with ‘unseen’ ones. Recent studies aim to leverage Vision-Language Models (e.g., CLIP [12]) to develop open-world 3D models that generalize beyond ‘seen’ 3D data, enabling zero-shot handling of various 3D tasks.
@@ 3.2 Depth-Aligned Image Generation
A total of 524,700 depth-aligned images are generated from ShapeNet [21], a leading public 3D CAD dataset containing approximately 52,470 models, each annotated with semantic metadata. To generate these images, a point cloud of 10,000 points is sampled from each CAD model, aligning with prior experimental protocols [18, 17]. For each point cloud, 10 views of depth maps are projected using the proposed contour-aware projection method. Subsequently, the ControlNet v1.1 [40] depth model produces depth-aligned images for each contour-aware depth map view, using the CAD model’s metadata as text input and the inverse of the depth map $1-D$ as image generation control. This approach ensures that the generated images remain consistent with the depth maps both geometrically and semantically, while also adding texture diversity across different views. ControlNet utilizes $1-D$ instead of $D$ for controlling image outputs because it is predominantly pre-trained on depth images where brighter regions indicate closer proximity. Please refer to Appendix A.2 for the positive and negative prompts used in ControlNet to achieve high-fidelity depth-aligned image generation.
@@ A.2 Implementation Details
A.3 New Asset
@@
A.4 Additional Quantitative Analysis
@@
Figure 5: Effect of the number of views on OpenDlign’s zero-shot performance on the ScanObjectNN Dataset [47]
@@
A.5 Additional Qualitative Analysis
@@
More Visualizations of Cross-Modal Retrieval. Additional examples of 3D shape retrieval that showcase the cross-modal retrieval capabilities of OpenDlign are illustrated in Fig. 7. The visualization results distinctly demonstrate that the 3D shapes retrieved are in semantic alignment with the input image and text queries.
@@
Examples of Low-Quality Depth-Aligned Images. Fig. 8 showcases examples of poor-quality depth-aligned images. For instance, the image diffusion model may interpret ‘monitor’ as a person observing something, or it might create unrealistic images such as a washing machine or a table made of stone or grass. Additionally, it might depict a tent floating in the sky, which deviates from real-world expectations. These examples underscore the need to eliminate inferior depth-aligned images for more generalized multimodal alignment.

### v3

# OpenDlign: Open-World Point Cloud Understanding with Depth-Aligned Images

@@ Abstract
Recent open-world 3D representation learning methods using Vision-Language Models (VLMs) to align 3D point clouds with image-text information have shown superior 3D zero-shot performance. However, CAD-rendered images for this alignment often lack realism and texture variation, compromising alignment robustness. Moreover, the volume discrepancy between 3D and 2D pretraining datasets highlights the need for effective strategies to transfer the representational abilities of VLMs to 3D learning. In this paper, we present OpenDlign, a novel open-world 3D model using depth-aligned images generated from a diffusion model for robust multimodal alignment. These images exhibit greater texture diversity than CAD renderings due to the stochastic nature of the diffusion model. By refining the depth map projection pipeline and designing depth-specific prompts, OpenDlign leverages rich knowledge in pre-trained VLM for 3D representation learning with streamlined fine-tuning. Our experiments show that OpenDlign achieves high zero-shot and few-shot performance on diverse 3D tasks, despite only fine-tuning 6 million parameters on a limited ShapeNet dataset. In zero-shot classification, OpenDlign surpasses previous models by 8.0% on ModelNet40 and 16.4% on OmniObject3D. Additionally, using depth-aligned images for multimodal alignment consistently enhances the performance of other state-of-the-art models.
@@ 1 Introduction
3D understanding, including tasks like 3D object classification [1, 2], 3D scene segmentation [3], and 3D reconstruction [4, 5], is becoming increasingly pivotal in real-world applications like augmented/virtual reality [6, 7], autonomous vehicles [8]. Traditional 3D models [1, 2, 9, 10, 11] are closed-world, recognizing only pre-defined categories and struggling with ‘unseen’ ones. Recent studies aim to leverage Vision-Language Models (e.g., CLIP [12]) to develop open-world 3D models that generalize beyond ‘seen’ 3D data, enabling zero-shot handling of various 3D tasks.
@@ 3.2 Depth-Aligned Image Generation
A total of 524,700 depth-aligned images are generated from ShapeNet [21], a leading public 3D CAD dataset containing approximately 52,470 models, each annotated with semantic metadata. To generate these images, a point cloud of 10,000 points is sampled from each CAD model, aligning with prior experimental protocols [18, 17]. For each point cloud, 10 views of depth maps are projected using the proposed contour-aware projection method. Subsequently, the ControlNet v1.1 [40] depth model produces depth-aligned images for each contour-aware depth map view, using the CAD model’s metadata as text input and the inverse of the depth map $\frac{1}{D}$ as image generation control. This approach ensures that the generated images remain consistent with the depth maps both geometrically and semantically, while also adding texture diversity across different views. ControlNet utilizes $\frac{1}{D}$ instead of $D$ for controlling image outputs because it is predominantly pre-trained on depth images where brighter regions indicate closer proximity. Please refer to Appendix A.2 for the positive and negative prompts used in ControlNet to achieve high-fidelity depth-aligned image generation.
@@ 5 Conclusion
6 Acknowledgement and Disclosure of Funding
We extend our gratitude to Maojun Zhang, Ranran Huang, Jiangnan Ye, Ziyang Chen, and Chengzu Li for their valuable discussions and insightful feedback on the early drafts of this work. This research was supported by the Imperial College President’s PhD Scholarships 2023.
@@ A.2 Implementation Details
A.3 Discussion
How to ensure multi-view geometry consistency in depth-aligned images? There is no geometric inconsistency because depth-aligned images are generated view by view, each paired with its respective depth map. ControlNet’s [40] conditional control ensures the generated images maintain the same shape and pose as the input depth map, while a text prompt preserves object identity. Since these depth maps originate from the same 3D point cloud, the resulting images remain geometrically consistent across views. Fig. 5 visually demonstrates the generated images are geometry-consistent.
Figure 5: Examples of multi-view depth maps and their corresponding depth-aligned images.
Why does increasing texture diversity in multi-view images positively impact open-world 3D learning? First, more effective knowledge transfer is achieved by leveraging inconsistent textures in multi-view generated images, allowing depth maps to align with diverse sets of RGB images, which enhances the transfer of rich 2D knowledge embedded in CLIP to 3D representation learning. Second, generating geometry-consistent but texture-inconsistent multi-view images benefits 3D representation learning by focusing on invariant features like object shape, size, and contour, which remain robust to texture variations. Third, similar to previous methods [15, 17, 18], OpenDlign pairs a single-view depth map with a single-view RGB image for contrastive learning, treating each view independently and eliminating the need for texture consistency across views. Lastly, texture features are generally unreliable in both 3D and 2D data, as shown by the common use of color jittering for data augmentation in 2D contrastive learning. Increasing texture diversity is a standard approach to improve the robustness of contrastive learning models.
A.4 New Asset
@@
A.5 Additional Quantitative Analysis
@@
More Results for Zero-Shot 3D Object Detection. We followed the setting in [52] to compare our method with two 3D open-world approaches, focusing on object detection tasks OV-3DET [52] and CoDA [53] on the Scannet [51] dataset, as shown in Table 9. The results demonstrate that OpenDlign performs comparably to the state-of-the-art CoDA method, which is promising given that OpenDlign is not specifically designed for open-world 3D detection.
Table 9: Comparison of 3D open-vocabulary object detection methods in the same setting as OV-3DET on ScanNet. ’Mean’ represents the average precision across all 20 categories.
Methods Mean toilet bed chair sofa dresser table cabinet bookshelf pillow sink
OV-3DET 18.02 57.29 42.26 27.06 31.50 8.21 14.17 2.98 5.56 23.00 31.60
CoDA 19.32 68.09 44.04 28.72 44.57 3.41 20.23 5.32 0.03 27.95 45.26
OpenDlign (Ours) 19.27 58.30 41.95 34.13 32.25 12.49 18.12 3.89 4.15 25.41 34.53
Methods bathtub refrigerator desk nightstand counter door curtain box lamp bag
OV-3DET 56.28 10.99 19.27 0.77 0.31 9.59 10.53 3.78 2.11 2.71
CoDA 50.51 6.55 12.42 15.15 0.68 7.95 0.01 2.94 0.51 2.02
OpenDlign (Ours) 58.20 8.74 17.85 5.49 1.53 11.46 6.82 5.03 1.54 3.17
@@
Figure 6: Effect of view count on OpenDlign’s zero-shot performance on ScanObjectNN [47]
@@
A.6 Additional Qualitative Analysis
@@
More Visualizations of Cross-Modal Retrieval. More examples of 3D shape retrieval that showcase the cross-modal retrieval capabilities of OpenDlign are illustrated in Fig. 8. The results demonstrate that the 3D shapes retrieved are in semantic alignment with the image and text queries.
@@
Examples of Low-Quality Depth-Aligned Images. Fig. 9 showcases bad examples of depth-aligned images. For instance, the diffusion model may interpret ‘monitor’ as a person observing something, or it might create unrealistic images such as a washing machine made of stone. Additionally, it might depict a tent floating in the sky, which deviates from real-world expectations. These examples underscore the need to eliminate inferior depth-aligned images for more generalized alignment.